# Dark Matter Particle Flux in a Dynamically Self-consistent Milky Way Model

### v1

@@ 1 Introduction
Taking these observations into account, one can infer the distribution of DM in the Milky Way galaxy and draw conclusions about its local density. This information can be used to estimate the magnitude of seasonal variations and directional effects, which could provide evidence for the discovery of DM (Spergel, 1988). Seasonal variation is expected as Earth passes through varying densities of DM, depending on its relative velocity to the galactic center. In June, Earth’s motion around the Sun aligns with the Solar System’s motion, while in December, it opposes it. This would lead to a signal modulation with an annual periodicity.
@@ 2 Galactic halo models
To estimate this modulation, a model of the local density and velocity distributions is needed. A fundamental framework is provided by the standard halo model (SHM) (Drukier et al., 1986), which posits that the Milky Way is embedded in a roughly spherical and isotropic DM halo extending beyond the visible components of the galaxy. However, recent observations from the Gaia space observatory have revealed more complexity in our galaxy’s halo structure, particularly with the discovery of the Gaia-Sausage-Enceladus (GSE) (Belokurov et al., 2018; Helmi et al., 2018) – an ancient merger between the Milky Way and a smaller dwarf galaxy, which likely occurred 8–10 Gyr ago and manifests itself as a population of halo stars with highly eccentric orbits.
In light of this discovery, Evans et al. (2019) introduced an updated DM model, referred to as SHM++, which includes a DM counterpart to the GSE. Both the main component (nearly identical to SHM) and the GSE are assumed to follow smooth, spherically-symmetric density profiles and have Gaussian velocity distributions truncated at the escape velocity $v_{\mathrm{esc}}$, but while the former is still isotropic, the latter is assumed to have the same high radial anisotropy as stars, i.e. the anisotropy coefficient $\beta\equiv 1-\sigma_{\mathrm{tan}}^{2}/(2\sigma_{\mathrm{rad}}^{2})\simeq 0.9$. The circular speed $v_{0}$, the escape speed, the local DM density, the fractional contribution of GSE and its anisotropy were updated based on the recent measurements, predominantly from Gaia. They then derived the analytical velocity distribution and other observables from this model.
However, the truncated Gaussian velocity distribution in the SHM++ model is not very realistic. Although the true distribution function (DF) $f({\boldsymbol{x}},{\boldsymbol{v}})$ of the DM halo is not observationally constrained, the Jeans theorem stipulates that in a steady state, it must be a function of the integrals of motion $\mathcal{I}(\boldsymbol{x},\boldsymbol{v};\;\Phi)$ in the potential $\Phi$. Recently, Binney & Vasiliev (2023) constructed a dynamically self-consistent model of the Milky Way constrained by the Gaia data, which provides a plausible choice for such a DF. In the present paper, we explore the implications of this model on the DF velocity distribution in the Solar neighbourhood.
@@ 3 Methods
and their parameters are given in table 1 (the normalization factor $A$ is computed from the prescribed total mass). The remaining parameters describing the stellar components are the same as in the original paper. In what follows, we refer to updated Milky Way model with a two-component DM halo as the self-consistent Anisotropic Halo Model (scAHM).
Figure 1 shows the radial profiles of density and three components of velocity dispersion along the $x$ axis for both DM halo components: main halo by solid lines and GSE by dashed lines. The density profiles are designed to be similar at all radii, with the GSE component being 4$\times$ less massive. The velocity dispersion of the main halo is nearly isotropic, while GSE has $\sigma_{r}$ much higher than both tangential velocity components.
@@
with nucleus mass $m_{N}$ and its reduced mass with the DM particle $\mu_{N}$. The term $\sigma_{T}$ refers to the DM nucleus scattering cross-section
@@ 4.1 Galactocentric velocity distributions
The velocity distribution of the main halo component of scAHM slightly differs from the Gaussian of the SHM, having a steeper decline and a smooth transition to zero at $v>v_{\rm esc}$ instead of a sharp truncation. On the other hand, the GSE component has a significantly non-Gaussian distribution in all three components. The radial velocity distribution is much wider than a Gaussian and has a distinctly top-hat shape, while still smoothly transitioning to zero at the escape velocity. By contrast, the distributions of azimuthal and to a lesser extent vertical velocities have a narrower central peak than a Gaussian, and therefore a much smaller dispersion than the radial component. This is a natural consequence of the imposed radial anisotropy of this component: most of kinetic energy is contained in the radial motion, thus the other two components are narrowly peaked around zero. The shape the velocity distribution is the most important difference between our model and SHM++, which also includes a radially-anisotropic GSE component, but retains a truncated Gaussian approximation to its velocity distribution.
@@ 4.3 Expected Event Rates
As target nuclei, we chose Silicium, Argon, Germanium, and Xenon in a hypothetical experiment. For simplicity, we assumed the dark matter candidate to have a mass of $m_{\chi}=5$ GeV/c2. The plots for the WIMP candidates with higher masses can be found in the appendix.
@@ 5 Conclusion
A limitation of our model is the exclusion of potential DM halo perturbations from external galaxies, notably the Large Magellanic Cloud (LMC). Simulations (Smith-Orlik et al., 2023) suggest the LMC could induce a high-speed tail in the local DM velocity distribution. While a full consideration is beyond our scope, we justify this exclusion by assuming the GSE merger dominates the non-Gaussian features, with the LMC’s influence representing a second-order correction.
Altogether, the impact of the GSE component would be most clearly observable in the direction sensitive detectors, where its distinct velocity distribution features could be directly measured.
@@ Acknowledgments
We would like to thank Laura Baudis for her insightful comments and valuable suggestions, which greatly helped improve our paper.

### v2

@@ 1 Introduction
Taking these observations into account, one can infer the distribution of DM in the Milky Way galaxy and draw conclusions about its local density. This information can be used to estimate the magnitude of seasonal variations and directional effects, which could provide evidence for the discovery of DM (Spergel, 1988). Seasonal variation is expected as Earth passes through varying densities of DM, depending on its relative velocity to the galactic center. In June, Earth’s motion around the Sun aligns with the Solar System’s motion, while in December, it opposes it. This would lead to a signal modulation with annual periodicity.
@@ 2 Galactic halo models
To estimate this modulation, a model of the local density and velocity distributions is needed. A fundamental framework is provided by the standard halo model (SHM) (Drukier et al., 1986), which posits that the Milky Way is embedded in a roughly spherical and isotropic DM halo extending beyond the visible components of the galaxy. However, the SHM has become outdated in multiple ways, particularly in its estimate of the local dark matter density, which was originally set at 0.3 GeV/cm3 but has been revised upward to 0.4–0.5 GeV/cm3 (see figure 1 in de Salas & Widmark (2021) for a compilation of recent results). Recent observations from the Gaia space observatory have even revealed more complexity in our galaxy’s halo structure, particularly with the discovery of the Gaia-Sausage-Enceladus (GSE) (Belokurov et al., 2018; Helmi et al., 2018) – an ancient merger between the Milky Way and a smaller dwarf galaxy, which likely occurred 8–10 Gyr ago and manifests itself as a population of halo stars with highly eccentric orbits.
In light of this discovery, Evans et al. (2019) introduced an updated DM model, referred to as SHM++, which includes a DM counterpart to the GSE. Both the main component (nearly identical to SHM) and the GSE are assumed to follow smooth, spherically-symmetric density profiles and have Gaussian velocity distributions truncated at the escape velocity $v_{\mathrm{esc}}$, but while the former is still isotropic, the latter is assumed to have the same high radial anisotropy as stars, i.e. the anisotropy coefficient $\beta\equiv 1-\sigma_{\mathrm{tan}}^{2}/(2\sigma_{\mathrm{rad}}^{2})\simeq 0.9$. The circular speed $v_{0}$, the escape speed, the local DM density, the fractional contribution of GSE and its anisotropy were updated based on recent measurements, predominantly from Gaia. They then derived the analytical velocity distribution and other observables from this model.
However, the truncated Gaussian velocity distribution in the SHM++ model is not very realistic. Although the true distribution function (DF) $f({\boldsymbol{x}},{\boldsymbol{v}})$ of the DM halo is not observationally constrained, the Jeans theorem stipulates that in a steady state, it must be a function of the integrals of motion $\mathcal{I}(\boldsymbol{x},\boldsymbol{v};\;\Phi)$ in the potential $\Phi$. Recently, Binney & Vasiliev (2023) constructed a dynamically self-consistent model of the Milky Way constrained by Gaia data, which provides a plausible choice for such a DF. In the present paper, we explore the implications of this model on the DF velocity distribution in the Solar neighbourhood.
@@ 3 Methods
and their parameters are given in table 1 (the normalization factor $A$ is computed from the prescribed total mass). The remaining parameters describing the stellar components are the same as in the original paper. In what follows, we refer to this updated Milky Way model with a two-component DM halo as the self-consistent Anisotropic Halo Model (scAHM). We stress that while the DF of the GSE component needs to have strong radial anisotropy, there is no fundamental reason why it should follow the above functional form; here we chose it purely for convenience, as one of the standard models implemented in AGAMA. Ideally, one should look to cosmological simulations of Milky Way analogues with GSE-like merger history and find an appropriate form for the DF of the accreted component. Lane & Bovy (2025) analysed an ensemble of Milky Way analogues from the IllustrisTNG simulation and concluded that DFs with a constant or radially increasing anisotropy of the Osipkov–Merritt form (Cuddeford, 1991) can describe the merger debris well; however, they only examined the velocity dispersion profiles, not the specific functional form of these DFs.
Figure 1 shows the radial profiles of density and three components of velocity dispersion along the $x$ axis for both DM halo components: main halo by solid lines and GSE by dashed lines. The density profiles are designed to be similar at all radii, with the GSE component being 4$\times$ less massive. The velocity dispersion of the main halo is nearly isotropic, while GSE has $\sigma_{r}$ much higher than both tangential velocity components. However, the level of anisotropy in our model ($\beta\simeq 0.65$) is not as high as inferred from observations, which may be a limitation of our chosen DF family. Thus our conclusions about the effect of radial anisotropy on DM detection can only be regarded as qualitative.
@@
with nucleus mass $m_{N}$ and its reduced mass with the DM particle $\mu_{N}$. The factor $\sigma_{T}$ refers to the DM nucleus scattering cross-section
@@ 4.1 Galactocentric velocity distributions
The velocity distribution of the main halo component of scAHM slightly differs from the Gaussian of the SHM, having a steeper decline and a smooth transition to zero at $v>v_{\rm esc}$ instead of a sharp truncation. On the other hand, the GSE component has a significantly non-Gaussian distribution in all three components. The radial velocity distribution is much wider than a Gaussian and has a distinctly top-hat shape, while still smoothly transitioning to zero at the escape velocity. By contrast, the distributions of azimuthal and to a lesser extent vertical velocities have a narrower central peak than a Gaussian, and therefore a much smaller dispersion than the radial component. This is a natural consequence of the imposed radial anisotropy of this component: most of the kinetic energy is contained in the radial motion, thus the other two components are narrowly peaked around zero. The shape the velocity distribution is the most important difference between our model and SHM++, which also includes a radially-anisotropic GSE component, but retains a truncated Gaussian approximation to its velocity distribution. We note that the narrow spike in $f(v_{\phi})$ (green curve) at zero might be an unphysical feature of our chosen DF family; however, in the case of a Cuddeford (1991) DF used by Lane & Bovy (2025), the spike is even more pronounced and manifests itself in both $v_{\phi}$ and $v_{z}$ components, although this hardly makes this DF any more realistic than our choice. Whatever the actual DF of the GSE component is, it likely has a peak (but not necessarily a spike) at small $v_{\phi}$ and $v_{z}$, perhaps lying somewhere between the red and green curves in our figure.
@@ 4.3 Expected Event Rates
As target nuclei, we chose Silicon, Argon, Germanium, and Xenon in a hypothetical experiment. For simplicity, we assumed the dark matter candidate to have a mass of $m_{\chi}=5$ GeV/c2. The plots for the WIMP candidates with higher masses can be found in the appendix.
@@ 5 Conclusion
Studying the effect of the GSE component on the differential event rates, we observe a slight decrease in differential event rates at high recoil energies, which is counterbalanced by a modest increase in overall differential event rates due to changes in other parameters.
@@
A limitation of our model is the exclusion of potential DM halo perturbations from external galaxies, notably the Large Magellanic Cloud (LMC). Simulations (e.g., Besla et al., 2019; Donaldson et al., 2022; Smith-Orlik et al., 2023) suggest that the LMC could induce a high-speed tail in the local DM velocity distribution. While a full consideration is beyond our scope, we justify this exclusion by assuming that the GSE merger dominates the non-Gaussian features, with the LMC’s influence representing a second-order correction.
Altogether, the impact of the GSE component would be most clearly observable in direction-sensitive detectors, where its distinct velocity distribution features could be directly measured.
@@ Acknowledgments
We would like to thank Laura Baudis and the anonymous referee for their insightful comments and valuable suggestions, which greatly helped to improve our paper. EV acknowledges support from an STFC Ernest Rutherford fellowship (ST/X004066/1).